lucy makes a rectangle with the perimeter of 30. The lenght is twice the width. how do you work out the area?
Answer by ptaylor(2198)

OK. The area of a rectangle is length multiplied width or: A=LW
The perimeter of a rectangle is twice the length plus twice the width or P=2L+2W
Now in this problem, let W=the width of the rectangle
Then the length, L=2W
So the perimeter would be:P=30= 2(2W)+2W (since L=2W)
So, our equation to solve to get the length and width is:
30=2(2W)+2W get rid of parens
30=4W+2W collect like terms
30=6W divide each side by 6
W=5 units
Then L=2W=2*5=10 units
Now we can look at the area:
A=L*W=5*10=50 square units
CK
30=2*10+2*5
30=20+10
30=30
Hope this helps----ptaylor